20151209 UT DDT

Mountain Manager: John Little (on mountain)
OSA: David Gonzalez Huerta (on mountain)
Instrument Specialist: John Morris (on mountain)
Observers for DDT: Olga Kuhn, Barry Rothberg (LBTO) (in Tucson)
OSURC observers working before/after: Jonathan Brown (OSU), Evan Skillman (UM) (in Tucson)
ISA for OSURC observers: Olga Kuhn (LBTO) (in Tucson)

Transit observation.

Cirrus present, but transparency is pretty good.

03:47: Slewing to field.

03:50: dofpia. Needed z8=+2500 on DX. All aberrations were small except defocus a bit big... "S" to stop dofpia.

03:57: send copoint OB.

04:01: ask for Z go = -1.0 on SX and Z go = -1.3 on DX. Send test OB epic_v4_test to take a 100s on g and and 30s on z.
  • SX: max cnts 55000, diam = 15
  • DX: max cnts 9000, diam = 80

04:11: drop 60s:30s and Z=-1.2 on both
  • SX: (LBC Blue channel problem --- stuck in exposure. lbckill/lbcstart)
  • DX: 10,000 d=44

04:18: test with 60s:60s and Z=-1.2 on both
  • SX: -1.2 -> 10,000 on target
  • DX: -1.0 -> 20,000 on target

04:27: sent epic_v5 which uses 120s:60s on g:z and zgo = -1.2 SX and zgo=-1.0 DX
  • there is something strange about the global offsets, but we should have the appropriate count levels.
    • for reference, D=30 pixels on SX and D=78 pixels on DX
  • 43141 b, 43026 r have ~19,000/20,000 counts at peak.

05:35: stability looks good for both channels - counts and diameter of pupil.
  • sky looks clearer than when it was started.

05:50: Shutter open times: the filenames contain the UT when the shutter opened. This time comes from the instrument PCs and the times should be synced with the GPS.
  • On red - I see UT 05:52:40 for red image 055244 and I see UT 05:54:22 for file 055216.

06:41: The g images are much brighter than they were previously --- 49,000 max. But the red are the same. Paused and edit the OB to use 60s:60s for both g:z. Diameter of blue pupil may have dropped a bit and the illumination pattern changed. Started around 06:30, and by 063705 max value is ~49,000. The target crossed the meridian at 06:22 --- I think the change in pupil illumination is related to the meridian crossing.

06:45: Start OB with 60s:60s and no change to defocus. This helped, but still the pupil shape is changing as we round out this series of observations. On 065251, peak at >~46,000, on 065357, 51,374, but on 065628, 49,000.

07:01: done. Last exposures were B: 065814 and R: 065720. I had clicked pause to edit the scale factor, but it did not pause until much later, finally after these two. Did not resume since we were so close to the 07:00:00 end. The data sequence started about 30-min late - we stopped OSURCs queue about 15 min too late, and setup required about 45 total. We did not take in-focus images on either side, but the g images are close enough to focus to allow easy id of the field.

g flats were taken the night before last (1207) and biases were taken within the past 2 days. z flats will be tried this morning, weather permitting.

The cirrus again on the horizon according to the all-sky , but during the transit, conditions overhead seemed clear.

06:00 Reconfiguring left side MODS->LBCB. (LBCR is in the beam already).

06:11 Sending telescope to ra=10:24, dec=27 (lst=11:00). There is definitely cirrus. It shows up more along the horizons than near zenith on the all-sky, but it is probably all over the sky. See the view from MGIO base camp taken ~7am (14 UT) below (the mountains are not snow-covered! There must be something wrong with the camera's color scheme). I'll take flats, but do note they were taken in cirrus.
  • g and z use exptimes 1:1.3 to get equal counts in each channel (better than default 1:2.2 scaling)

06:30: start dithered set of 5 at PA=0 using scale = 3 so exptimes = 3:(3*1.3) for g:z
  • 14k-22k for z and 11k-19k for g.
  • repeat 26k-41k for z and 25-47 for g
  • at PA=180, start dithered set with scale=1.5, then 0.7, then 0.4 ---> g 41k,24,28k, and z 41,24,25,19
  • at PA=180, repeat with 0.4 scale
  • at PA=180, repeat z only with 0.4 scale

06:50 David is closing up.

06:56: Start a set of 25 biases. Dome is dark, telescope & enclosure stationary. First 18 are red-only, a remnant of disabling the blue channel when it got too bright for flats through these.
  • r 140421 is bad
  • set of red and blue biases started 14:07 UT.

-- %USERSIG{OlgaKuhn - 2015-12-09}%

Comments

 
  • sky conditions ~14:00 UT (7am) 20151209:
    cirrusy_sky20151209_14UT.jpg
I Attachment Action Size Date Who Comment
cirrusy_sky20151209_14UT.jpgjpg cirrusy_sky20151209_14UT.jpg manage 23 K 09 Dec 2015 - 13:59 OlgaKuhn sky conditions ~14:00 UT (7am) 20151209
Topic revision: r5 - 09 Dec 2015, OlgaKuhn
This site is powered by FoswikiCopyright © by the contributing authors. All material on this collaboration platform is the property of the contributing authors.
Ideas, requests, problems regarding Foswiki? Send feedback